# Constraints from Entanglement Wedge Nesting for Holography at a Finite Cutoff

### v2

@@ 2.2 Candidate RT Surfaces in AdS3 with an ETW Brane
There are two possible configurations of the extremal surfaces. Either, the extremal surface consists of two disconnected pieces which are separately homologous to $A$ and $B$. We will denote those surfaces by $\chi_{\text{dis}}(A)$, $\chi_{\text{dis}}(B)$, respectively, where the subscript stands for disconnected. Explicitly, $\chi_{\text{dis}}(A)$ is given by
@@ 2.3 General Classification of Extremal Surfaces in Poincare AdS3
In usual AdS/CFT, i.e. when the spacetime boundary is at asymptotic infinity, one only considers RT curves that have $|k|<1$ (i.e. circles and ellipses), since all the extremal surfaces have to end on the $z=0$ boundary. RT surfaces with $|k|\geq 1$ may start on the asymptotic boundary at $z=0$ but never return. However, in the presence of a cutoff, such as the brane under consideration, we can now attach one endpoint of the extremal surface to the brane for which generally $z>0$. This allows for novel configurations where the RT curve has $|k|\geq 1$.
@@ 2.4 Constructing the Entanglement Wedges
Once the candidate RT surfaces are known, we can construct the associated entanglement wedges. In standard holography, the entanglement wedge is defined by taking a boundary subregion $U$ together with its associated RT surface $\chi(U)$. One then considers a partial Cauchy surface, $\Sigma_{U}$, whose boundary is given by $\partial\Sigma_{U}=U\cup\chi(U)$. The entanglement wedge of the boundary subregion $U$, denoted $\mathcal{W}_{E}(U)$, is then given by the domain of dependence of the partial Cauchy slice $\Sigma_{U}$. It is tempting to suggest that to construct the entanglement wedge $\mathcal{W}_{E}(U)$ to an associated boundary region $U$ one can simply define an extended entanglement wedge using $\bar{\chi}(U)$ and then cut off the extended wedge with brane and identify the remaining portion in $\mathcal{M}_{\text{phys}}$ as $\mathcal{W}_{E}(U)$. As we shall see in our discussions, such a prescription is generally not valid in the cases where our boundary subregions lie on cutoff surfaces such as the brane.
@@
Figure 3: An example of $\mathcal{W}_{E}(A)$. The intersection of the past and future cones is the RT surface $\chi_{\text{dis}}(A)$ shown in green. The dotted “diamond” is the causal diamond on the boundary associated with the interval $A$ which horizontally splits the diamond to the future and past of the interval.
In our setup we also have spatial boundary subregions that (at least partially) live on the brane, namely $B$ and $A\cup B$. We already know that the RT curves associated with $B$ and $A\cup B$ are given by $\chi_{\text{dis}}(B)$ and $\chi_{\text{con}}(A\cup B)$ respectively. We define the entanglement wedge of the subregion $B$ denoted, $\mathcal{W}_{E}(B)$, as the domain of dependence of the partial Cauchy slice $\Sigma_{B}$ whose boundary is given by $\partial\Sigma_{B}=B\cup\chi_{\text{dis}}(B)$. In a similar manner we define the the entanglement wedge of subregion $A\cup B$, denoted $\mathcal{W}_{E}(A\cup B)$, as the domain of dependence of the partial Cauchy slice $\Sigma_{A\cup B}$ whose boundary is given by $\partial\Sigma_{A\cup B}=A\cup B\cup\chi_{\text{con}}(A\cup B)=A\cup B\cup\chi_{1}\cup\chi_{2}$, where we remind the reader that $\chi_{1}$ and $\chi_{2}$ are the extremal surfaces which connect the boundaries of $A$ and $B$.
@@ 2.4.1 Spacelike Separated Points from RT Curves in Poincare AdS3
Suppose we have an extremal line segment, $\chi$, parameterized in terms of the Poincare coordinate $x\in[x_{L},x_{R}]$ (i.e. we have a curve $z=z_{RT}(x)$ and $t=t_{RT}(x)$). At the endpoints of the line segment we introduce planes $\sigma_{L,R}$ that are orthogonal to the tangent vectors,
@@
Importantly, in general, if we have an extremal surface $\bar{\chi}$ that lives in AdS3 and another extremal surface $\chi=\bar{\chi}\cap\mathcal{M}_{\text{phys}}$, the set of points in $\mathcal{M}_{\text{phys}}$ which are spacelike separated from $\chi$ will generally differ from the set of spacelike separated points from $\bar{\chi}$ due to the finite extent of the extremal surfaces.
In particular, one could naively extended all extremal surfaces behind the brane, $\chi\to\bar{\chi}$, and defined spacelike separation in the physical spacetime $\mathcal{M}_{\text{phys}}$ as the restriction of spacelike separation in the extended spacetime to points in $\mathcal{M}_{\text{phys}}$. As we will see in Section 2.4.2, our approach and this naive approach will correctly identify $\mathcal{W}_{E}(B)$. However, when constructing $\mathcal{W}_{E}(A\cup B)$ in the connected phase in Section 2.4.3 the naive approach will not correctly give the entanglement wedge. It is also worth mentioning that if one defines an extended wedge $\bar{\mathcal{W}}$ in the extended spacetime using $\bar{\chi}$ we know by construction that null sheet enclosing $\bar{\mathcal{W}}$ will always fully enclose a co-dimension 0 region between $\bar{\chi}\cap\mathcal{M}_{\text{phys}}$ and the brane. However, it is possible for portions of the null sheet in front of the brane to be generated by geodesics that originate from $\bar{\chi}$ behind the brane. Our prescription relies only on null rays emitted from $\bar{\chi}\cap\mathcal{M}_{\text{phys}}$. In fact, we can say that the entanglement wedge can be correctly identified using the naive prescription iff the null sheet which encloses $\bar{\mathcal{W}}\cap\mathcal{M}_{\text{phys}}$ can be completely generated by null congruences emitted from $\bar{\chi}\cap\mathcal{M}_{\text{phys}}$. Such is the case for $\mathcal{W}_{E}(B)$ but not for $\mathcal{W}_{E}(A\cup B)$.
@@ 2.4.2 Construction of $\mathcal{W}_{E}(B)$
Making use of the notation discussed immediately after Eq. (16) we can define and describe $\mathcal{R}_{L,M,R}$. In particular, $\mathcal{R}_{M}=\mathcal{V}_{+}(\sigma_{L})\cap\mathcal{V}_{-}(\sigma_{R})$ can be intuitively understood as the region wedged between $\sigma_{R}$ and $\sigma_{L}$. The region $\mathcal{R}_{L}=\mathcal{V}_{-}(\sigma_{L})$ can be understood as being between the planes $x=-\infty$ and $\sigma_{L}$ (i.e. to the “left” of $\mathcal{R}_{M}$). In a similar manner $\mathcal{R}_{R}=\mathcal{V}_{+}(\sigma_{R})$ is the region between the planes $x=+\infty$ and $\sigma_{R}$ (i.e. to the “right” of $\mathcal{R}_{M}$). We adopt the notation $SL_{Q}(\chi_{\text{dis}}(B))$ to denote the set of spacelike separated points to $\chi_{\text{dis}}(B)$ in the subregion $Q=\{L,M,R\}$. Using this notation we can write [^8] We obtain the definition of $SL_{M}(\chi_{\text{dis}}(B))$ from Eq. (LABEL:SLTowardsBdry) and Eq. (LABEL:SLAwayBdry) by setting $x_{\pm}=\frac{b_{1}+b_{2}}{2\sin\theta_{0}}$ and $t_{\pm}=t_{b}\pm R_{b}$. The “$+$” corresponds to the spacelike points towards the boundary and the “--” corresponds to the set of spacelike points expanding away from the boundary.,
@@ 2.4.3 Construction of $\mathcal{W}_{E}(A\cup B)$
Now that we have identified the entanglement wedges associated with the boundary interval $A$ and brane interval $B$ we can consider the construction of the entanglement wedge associated with $A\cup B$. In this case there are two possible phases that can exist for the entanglement wedge of $A\cup B$. The disconnected phase, which is given by the union of the entanglement wedges discussed thus far (i.e. $\mathcal{W}_{E}(A\cup B)=\mathcal{W}_{E}(A)\cup\mathcal{W}_{E}(B)$). However, in the connected phase it is no longer true that the entanglement wedge is the union of the disjoint wedges. To determine $\mathcal{W}_{E}(A\cup B)$ in the connected phase we will adopt the procedure of finding the entanglement wedge through the prescription of finding the spacelike region away from the boundary from $\chi_{1}$ and the spacelike region toward the boundary for $\chi_{2}$ and then taking the intersection of those regions and restricting to $\mathcal{M}_{\text{phys}}$. Roughly speaking, the region we will define will look like a “tube” that connects $A$ and $B$ regions on the brane-boundary system.
For each line segment we must begin by computing the relevant orthogonal planes. To begin, we note that both line segments $\chi_{1}$ and $\chi_{2}$ have their left endpoint on the conformal boundary so the planes for these points coincide for the conformal boundary and the region $\mathcal{R}_{L}=\varnothing$ [^9] As we showed in Appendix A.2 the plane “$\sigma_{L}$” coincides with the boundary which implies that $\mathcal{R}_{L}=\varnothing$.. We only need to worry about $\sigma_{R}$ for $\chi_{1,2}$.
The orthogonal plane to $\chi_{i}$ at the the point where $\chi_{i}$ intersects the brane is denoted $\sigma_{R,i}$, where $i=1,2$. The points on the plane satisfy
@@
Now that we have illustrated how we can find the set of spacelike regions from from $\chi_{1,2}$ we can use this to construct the $\mathcal{W}_{E}(A\cup B)$ in the connected phase. We start by considering $SL^{+}_{AdS_{3}}(\chi_{1})\cap SL^{-}_{AdS_{3}}(\chi_{2})$ where $SL^{+}_{AdS_{3}}(\chi_{1})$ is the set of spacelike points from $\chi_{1}$ expanding “outward” towards $\chi_{2}$ and $SL^{-}_{AdS_{3}}(\chi_{2})$ is the set of spacelike points from $\chi_{2}$ expanding “inward” towards $\chi_{1}$. The intersection of the regions in $\mathcal{M}_{\text{phys}}$ will trace a smooth “tube-shaped” codimension 0 region enclosed by null sheets emitted from $\chi_{1,2}$ as long as $\chi_{1}$ and $\chi_{2}$ are spacelike separated from each other [^10] Unlike in standard holography, when we have a cutoff $\chi_{1}$ is not generally guaranteed to be spacelike separated from $\chi_{2}$ whenever $A$ and $B$ are spacelike separated. This fact will reappear in our discussions of EWN in Section 3.3.. Furthermore, we can postulate the existence of a specific partial Cauchy slice $\Sigma_{A\cup B}$ with boundary $\partial\Sigma_{A\cup B}=A\cup B\cup\chi_{1}\cup\chi_{2}$ and we can identify the domain of dependence of $\Sigma_{A\cup B}$ with the set of points $SL^{+}_{AdS_{3}}(\chi_{1})\cap SL^{-}_{AdS_{3}}(\chi_{2})\cap\mathcal{M}_{\text{phys}}$. For this reason we identify
@@
Figure 9: This plot shows the intersection of $\mathcal{W}_{E}(B)$ with the brane which is the region enclosed by the solid black line. The intersection of the connected wedge $\mathcal{W}_{E}(A\cup B)$ with the brane is enclosed by the solid red line. The region enclosed by the dotted blue lines is the naive causal diamond that one would associate with the interval $B$ just based on the induced metric on the brane. In general the intersection of the entanglement wedges with the brane are not the naive causal diamond, but rather a proper subset of the naive causal diamond.
@@ 3 Entanglement Wedge Nesting in Poincaré AdS
In this section, we will study conditions on the relative location of constant time intervals $A$ and $B$, which are located on the asymptotic boundary and on the cutoff brane respectively, such that EWN is satisfied, that is, $\mathcal{W}_{E}(A),\mathcal{W}_{E}(B)\subseteq\mathcal{W}_{E}(A\cup B)$. Clearly, this condition is only non-trivial if $\mathcal{W}_{E}(A\cup B)\neq\mathcal{W}_{E}(A)\cup\mathcal{W}_{E}(B)$, ie., whenever we are in the connected phase. Assuming the latter, in Section 3.1 we derive a condition, presented in Eq. (43), on the time separation between $A$ and $B$ such that EWN holds. In Section 3.2 we study when the connected phase dominates and our conditions for EWN should be applied. We find that there are nontrivial configurations in which the RT surfaces are in the connected phase and EWN fails, although the intervals $A$ and $B$ are spacelike separated with respect to both the bulk and boundary metric. In Section 3.3, we argue that these violations appear whenever the RT surfaces in the connected or disconnected phase are not spacelike separated. In Section 3.4, we discuss how insisting that $A$ and $B$ be be spacelike separated in the bulk is generally not enough to ensure EWN will be satisfied. We connect this to the idea that the holographic theory on the cutoff surface is non-local which leads to a stronger requirement than spacelike separation of $A$ and $B$ to define independent subregions in non-local holographic theories. Finally, in Section 3.5, we compare our prescription for defining entanglement wedges/RT surfaces in cutoff holography with the restricted maximin prescription of Grado-White:2020wlb . Our results imply that while restricted maximin is designed such that EWN is satisfied under the condition that $A$ and $B$ are spacelike separated, the results obtained from restricted maximin disagree with a naive application of the RT formula if the “naive” RT surfaces corresponding to $A$ and $B$ are in causal contact. Even in the simplest cases, like the one we discuss, this can lead to RT surfaces which significantly differ from the naive expectation.

### v3

@@ 2.2 Candidate RT Surfaces in AdS3 with an ETW Brane
There are two possible configurations of the extremal surfaces. Either, the extremal surface consists of two disconnected pieces which are separately homologous to $A$ and $B$. We will denote those surfaces by $\chi_{\text{dis}}(A)$ and $\chi_{\text{dis}}(B)$, respectively, where the subscript stands for disconnected. Explicitly, $\chi_{\text{dis}}(A)$ is given by
@@ 2.3 General Classification of Extremal Surfaces in Poincare AdS3
In usual AdS/CFT, i.e. when the spacetime boundary is at asymptotic infinity, one only considers RT curves that have $|k|<1$ (i.e. circles and ellipses), since all the extremal surfaces have to end on the $z=0$ boundary. RT surfaces with $|k|\geq 1$ may start on the asymptotic boundary at $z=0$ but never return. [^7] At least in the case where we restrict ourselves to a single Poincare patch as we will be doing in this paper. However, in the presence of a cutoff, such as the brane under consideration, we can now attach one endpoint of the extremal surface to the brane for which generally $z>0$. This allows for novel configurations where the RT curve has $|k|\geq 1$.
@@ 2.4 Constructing the Entanglement Wedges
Once the candidate RT surfaces are known, we can construct the associated entanglement wedges. In standard holography, the entanglement wedge is defined by taking a boundary subregion $U$ together with its associated RT surface $\chi(U)$. One then considers a partial Cauchy surface, $\Sigma_{U}$, whose boundary is given by $\partial\Sigma_{U}=U\cup\chi(U)$. The entanglement wedge of the boundary subregion $U$, denoted $\mathcal{W}_{E}(U)$, is then given by the domain of dependence of the partial Cauchy slice $\Sigma_{U}$. It is tempting to suggest that to construct the entanglement wedge $\mathcal{W}_{E}(U)$ to an associated boundary region $U$ one can simply define an extended entanglement wedge using $\bar{\chi}(U)$ and then cut off the extended wedge with brane and identify the remaining portion in $\mathcal{M}_{\text{phys}}$ as $\mathcal{W}_{E}(U)$. As we shall see in our discussions, such a prescription is generally not valid in cases where our boundary subregions lie on cutoff surfaces such as the brane.
@@
Figure 3: An example of $\mathcal{W}_{E}(A)$. The intersection of the past cone and future cone is the RT surface $\chi_{\text{dis}}(A)$ shown in green. The dotted “diamond” is the causal diamond on the boundary associated with the interval $A$ which horizontally splits the diamond to the future and past of the interval.
In our setup we also have spatial boundary subregions that (at least partially) live on the brane, namely $B$ and $A\cup B$. We already know that the RT curves associated with $B$ and $A\cup B$ are given by $\chi_{\text{dis}}(B)$ and $\chi_{\text{con}}(A\cup B)$ respectively. We define the entanglement wedge of the subregion $B$ denoted, $\mathcal{W}_{E}(B)$, as the domain of dependence of the partial Cauchy slice $\Sigma_{B}$ whose boundary is given by $\partial\Sigma_{B}=B\cup\chi_{\text{dis}}(B)$. In a similar manner we define the entanglement wedge of subregion $A\cup B$, denoted $\mathcal{W}_{E}(A\cup B)$, as the domain of dependence of the partial Cauchy slice $\Sigma_{A\cup B}$ whose boundary is given by $\partial\Sigma_{A\cup B}=A\cup B\cup\chi_{\text{con}}(A\cup B)=A\cup B\cup\chi_{1}\cup\chi_{2}$, where we remind the reader that $\chi_{1}$ and $\chi_{2}$ are the extremal surfaces which connect the boundaries of $A$ and $B$.
@@ 2.4.1 Spacelike Separated Points from RT Curves in Poincare AdS3
Suppose we have an extremal line segment, $\chi$, parameterized in terms of the Poincare coordinate $x\in[x_{L},x_{R}]$ (i.e. we have a curve $z=z_{RT}(x)$ and $t=t_{RT}(x)$). At the endpoints of the line segment we introduce planes $\sigma_{L,R}$ that are orthogonal to the tangent vectors of the line segment,
@@
Importantly, in general, if we have an extremal surface $\bar{\chi}$ that lives in AdS3 and another extremal surface $\chi=\bar{\chi}\cap\mathcal{M}_{\text{phys}}$, the set of points in $\mathcal{M}_{\text{phys}}$ which are spacelike separated from $\chi$ will generally differ from the set of spacelike separated points from $\bar{\chi}$ due to the finite extent of the extremal surfaces. In certain cases it is possible to correctly identify the entanglement wedge by naively defining a larger “entanglement wedge” using $\bar{\chi}$ and then simply take the portion of the larger wedge which lies in $\mathcal{M}_{\text{phys}}$. However in general such a naive approach can fail.
As we will see in Section 2.4.2, our approach of using $\chi$ and this naive approach which uses $\bar{\chi}$ will correctly identify $\mathcal{W}_{E}(B)$. However, when constructing $\mathcal{W}_{E}(A\cup B)$ in the connected phase in Section 2.4.3 the naive approach will not correctly give the entanglement wedge. It is also worth mentioning that if one defines an extended wedge $\bar{\mathcal{W}}$ in the extended spacetime using $\bar{\chi}$ we know by construction that null sheet enclosing $\bar{\mathcal{W}}$ will always fully enclose a co-dimension 0 region between $\bar{\chi}\cap\mathcal{M}_{\text{phys}}$ and the brane. However, it is possible for portions of the null sheet in front of the brane to be generated by geodesics that originate from $\bar{\chi}$ behind the brane. Our prescription relies only on null rays emitted from $\bar{\chi}\cap\mathcal{M}_{\text{phys}}$. In fact, we can say that the entanglement wedge can be correctly identified using the naive prescription iff the null sheet which encloses $\bar{\mathcal{W}}\cap\mathcal{M}_{\text{phys}}$ can be completely generated by null congruences emitted from $\bar{\chi}\cap\mathcal{M}_{\text{phys}}$. Such is the case for $\mathcal{W}_{E}(B)$ but not for $\mathcal{W}_{E}(A\cup B)$.
@@ 2.4.2 Construction of $\mathcal{W}_{E}(B)$
Making use of the notation discussed immediately after Eq. (16) we can define and describe $\mathcal{R}_{L,M,R}$. In particular, $\mathcal{R}_{M}=\mathcal{V}_{+}(\sigma_{L})\cap\mathcal{V}_{-}(\sigma_{R})$ can be intuitively understood as the region wedged between $\sigma_{R}$ and $\sigma_{L}$. The region $\mathcal{R}_{L}=\mathcal{V}_{-}(\sigma_{L})$ can be understood as being between the planes $x=-\infty$ and $\sigma_{L}$ (i.e. to the “left” of $\mathcal{R}_{M}$). In a similar manner $\mathcal{R}_{R}=\mathcal{V}_{+}(\sigma_{R})$ is the region between the planes $x=+\infty$ and $\sigma_{R}$ (i.e. to the “right” of $\mathcal{R}_{M}$). We adopt the notation $SL_{Q}(\chi_{\text{dis}}(B))$ to denote the set of spacelike separated points to $\chi_{\text{dis}}(B)$ in the subregion $Q=\{L,M,R\}$. Using this notation we can write, [^9] We obtain the definition of $SL_{M}(\chi_{\text{dis}}(B))$ from Eq. (LABEL:SLTowardsBdry) and Eq. (LABEL:SLAwayBdry) by setting $x_{\pm}=\frac{b_{1}+b_{2}}{2\sin\theta_{0}}$ and $t_{\pm}=t_{b}\pm R_{b}$. The “$+$” corresponds to the spacelike points towards the boundary and the “--” corresponds to the set of spacelike points expanding away from the boundary.
@@ 2.4.3 Construction of $\mathcal{W}_{E}(A\cup B)$
Now that we have identified the entanglement wedges associated with the boundary interval $A$ and brane interval $B$ we can consider the construction of the entanglement wedge associated with $A\cup B$. In this case there are two possible phases that can exist for the entanglement wedge of $A\cup B$. The disconnected phase, which is given by the union of the entanglement wedges discussed thus far (i.e. $\mathcal{W}_{E}(A\cup B)=\mathcal{W}_{E}(A)\cup\mathcal{W}_{E}(B)$) and the connected phase. In the connected phase it is no longer true that the entanglement wedge is the union of the disjoint wedges(i.e. $\mathcal{W}_{E}(A\cup B)\neq\mathcal{W}_{E}(A)\cup\mathcal{W}_{E}(B)$). To determine $\mathcal{W}_{E}(A\cup B)$ in the connected phase we will adopt the procedure of finding the entanglement wedge through the prescription of finding the spacelike region away from the boundary from $\chi_{1}$ and the spacelike region toward the boundary for $\chi_{2}$ and then taking the intersection of those regions and restricting to $\mathcal{M}_{\text{phys}}$. Roughly speaking, the region we will define will look like a “tube” that connects $A$ and $B$ regions on the brane-boundary system.
For each line segment we must begin by computing the relevant orthogonal planes. To begin, we note that both line segments $\chi_{1}$ and $\chi_{2}$ have their left endpoint on the conformal boundary so the planes for these points coincide for the conformal boundary and the region $\mathcal{R}_{L}=\varnothing$. [^10] As we showed in Appendix A.2 the plane “$\sigma_{L}$” coincides with the boundary which implies that $\mathcal{R}_{L}=\varnothing$. We only need to worry about $\sigma_{R}$ for $\chi_{1,2}$.
The orthogonal plane to $\chi_{i}$ at the point where $\chi_{i}$ intersects the brane is denoted $\sigma_{R,i}$, where $i=1,2$. The points on the plane satisfy
@@
Now that we have illustrated how we can find the set of spacelike regions from $\chi_{1,2}$ we can use this to construct the $\mathcal{W}_{E}(A\cup B)$ in the connected phase. We start by considering $SL^{+}_{AdS_{3}}(\chi_{1})\cap SL^{-}_{AdS_{3}}(\chi_{2})$ where $SL^{+}_{AdS_{3}}(\chi_{1})$ is the set of spacelike points from $\chi_{1}$ expanding “outward” towards $\chi_{2}$ and $SL^{-}_{AdS_{3}}(\chi_{2})$ is the set of spacelike points from $\chi_{2}$ expanding “inward” towards $\chi_{1}$. The intersection of the regions in $\mathcal{M}_{\text{phys}}$ will trace a smooth “tube-shaped” codimension 0 region enclosed by null sheets emitted from $\chi_{1,2}$ as long as $\chi_{1}$ and $\chi_{2}$ are spacelike separated from each other. [^11] Unlike in standard holography, when we have a cutoff $\chi_{1}$ is not generally guaranteed to be spacelike separated from $\chi_{2}$ whenever $A$ and $B$ are spacelike separated. This fact will reappear in our discussions of EWN in Section 3.3. Furthermore, we can postulate the existence of a specific partial Cauchy slice $\Sigma_{A\cup B}$ with boundary $\partial\Sigma_{A\cup B}=A\cup B\cup\chi_{1}\cup\chi_{2}$ and we can identify the domain of dependence of $\Sigma_{A\cup B}$ with the set of points $SL^{+}_{AdS_{3}}(\chi_{1})\cap SL^{-}_{AdS_{3}}(\chi_{2})\cap\mathcal{M}_{\text{phys}}$. For this reason we identify
@@
Figure 9: This plot (which uses same parameters described in Figure 7) shows the intersection of $\mathcal{W}_{E}(B)$ with the brane which is the region enclosed by the solid black line. The intersection of the connected wedge $\mathcal{W}_{E}(A\cup B)$ with the brane is enclosed by the solid red line. The region enclosed by the dotted blue lines is the naive causal diamond that one would associate with the interval $B$ just based on the induced metric on the brane. In general the intersection of the entanglement wedges with the brane are not the naive causal diamond, but rather a proper subset of the naive causal diamond.
@@ 3 Entanglement Wedge Nesting in Poincaré AdS
In this section, we will study conditions on the relative location of constant time intervals $A$ and $B$, which are located on the asymptotic boundary and on the cutoff brane respectively, such that EWN is satisfied, that is, $\mathcal{W}_{E}(A),\mathcal{W}_{E}(B)\subseteq\mathcal{W}_{E}(A\cup B)$. Clearly, this condition is only non-trivial if $\mathcal{W}_{E}(A\cup B)\neq\mathcal{W}_{E}(A)\cup\mathcal{W}_{E}(B)$, ie., whenever we are in the connected phase. Assuming the latter, in Section 3.1 we derive a condition, presented in Eq. (43), on the time separation between $A$ and $B$ such that EWN holds. In Section 3.2 we study when the connected phase dominates and our conditions for EWN should be applied. We find that there are nontrivial configurations in which the RT surfaces are in the connected phase and EWN fails, even though the intervals $A$ and $B$ are spacelike separated with respect to both the bulk and boundary metric. In Section 3.3, we argue that these violations appear whenever the RT surfaces in the connected or disconnected phase are not spacelike separated. In Section 3.4, we discuss how insisting that $A$ and $B$ be spacelike separated in the bulk is generally not enough to ensure EWN will be satisfied. We connect this to the idea that the holographic theory on the cutoff surface is non-local which leads to a stronger requirement than spacelike separation of $A$ and $B$ to define independent subregions in non-local holographic theories. Finally, in Section 3.5, we compare our prescription for defining entanglement wedges/RT surfaces in cutoff holography with the restricted maximin prescription of Grado-White:2020wlb . Our results demonstrate that while restricted maximin is designed such that EWN is satisfied under the condition that $A$ and $B$ are spacelike separated, the results obtained from restricted maximin disagree with a naive application of the RT formula if the “naive” RT surfaces corresponding to $A$ and $B$ are in causal contact. Even in the simplest cases, like the one we discuss, this can lead to RT surfaces which significantly differ from the naive expectation.